OpenStudy (anonymous):

For 2x – 3y = –10, when x = 4, What is the value of y?

OpenStudy (amistre64):

when x=4; that only leaves one option for y

OpenStudy (anonymous):

put the value of x=4 in the given equ. 2x4-3y=-10 this implies 8-10=-3y -2=-3y and finally y=2/3

OpenStudy (anonymous):

when x=4 2(4)-3y=6 8-3y=6 8-6=3y 2=3y y=2/3